xenon detector status l.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
Xenon Detector Status PowerPoint Presentation
Download Presentation
Xenon Detector Status

Loading in 2 Seconds...

play fullscreen
1 / 49
matty

Xenon Detector Status - PowerPoint PPT Presentation

130 Views
Download Presentation
Xenon Detector Status
An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author. While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Xenon Detector Status Liquid Xenon Group

  2. Outline • Detector Setup • Operation • Performance • Problems and solutions

  3. Detector Setup

  4. PMT installation completed • All PMT successfully installed in Aug 2007 together with • LEDs • Alpha sources (plates and wires) • PT100 temperature sensors • Laser fibers • Surface level meter • < 200~300um gap btw inner slab and wall xenon g

  5. Sensors, LEDs, and alpha sources Capacitor-type surface level meter Pt100 Temp sensor 22 sensors in LXe 21 sensors on the wall • Two type of a sources • Plate 20 pieces • Wire 5x5 wires LED 3 different attenuation x 10

  6. Cabling and Al filler installation • Filler volume • Lateral 7.5lx6x2 + 4.5lx2x2 = 108l (design) 108 –7.5(US) – 4.5(DS) = 96l (reality) • Bottom 49l (design) 49 – 10 = 39l (reality) 7.5l 4.5l

  7. We were ready to close but… • The shape of the newly delivered sealing (metal gasket) was wrong ! • Does not fit the groove?! • Radii were correct but they made a mistake in bending • Sent back to Garlok and repaired • While waiting repair work, old gaskets were recycled with indium foil around the corner • Vacuum test OK • But leak of the order of 10-4 mbarl/sec with ~1 bar xenon in the cold vessel ~1 bar xenon outgas vacuum Warm vessel vacuum Xe leak Mass spectrometer

  8. RGA of leaking xenon • Doubly charged 132Xe (Xe++) is used to evaluate amount of xenon in the warm vessel • Production cross section with 110keV electron is ~40% of Xe+ production • Used as reference data for later use 10-4 mbarl/sec leak

  9. Closing the covers • Gaskets replaced with newly delivered ones on 5/Sep • Super-insulator on the cold vessel • 240 Nmm torque

  10. Pressure/leak test • Xenon gas (liquid) in the cold vessel • Mass spectrometer on the warm vessel • RGA data with recycled gasket used as a reference • Doubly charged 132Xe (Xe++) is used to evaluate amount of xenon in the warm vessel • Production cross section with 110keV electron is ~40% of Xe+ production 10-4 mbarl/sec leak

  11. Estimation of xenon leak rate Recycled gasket 10-4 mbarl/sec leak Room temperature xenon gas New gasket Room temperature xenon gas Better at low temperature

  12. Operation

  13. Gas/liquid system Gas-phase purifier Gas line Liquid line High pressure Storage Detector 1000L dewar Liquid-phase purifier

  14. Evacuation and liquefaction • Evacuation started on 5/Sep • Thanks to cryo pump (AISIN) • 6.9x10-3 Pa in 3 days • Cooling started on 10/Sep • Liquefaction started on 15/Sep • Surface level was monitored with • Temperature sensors • Level meter (long and short) Cryo pump Xenon gas 0.11 MPa 0.133 MPa

  15. Xenon liquid 0.25 MPa 0.11 MPa Liquid transfer • Liquid transfer by pressure difference between two cryostats through a vacuum insulated pipe • Started on 17/Sep and completed on 20/Sep • 10 liter/hour transfer speed • Xenon filling was done in 15 days after starting evacuation ~3m

  16. End of xenon filling • Additional 10 liter was transferred for assurance temperature Level meter

  17. How many liters of liquid xenon? • Cryostat volume 1200 liter • PMT and support 142 + 43 = 185 liter • Filler 135 liter • Lateral 96 liter • Bottom 39 liter • Cable 20~30 liter • Additionally transferred amount of xenon ~10 liter • 1200 – 185 – 135 – 25 + 10 = 865 liter • Consistent with remaining amount of xenon in the dewar

  18. Labview Slow Control megon00 PC in barrack MEG Central DAQ System (SCFE) For shift crew use Labview XEC PC 2 XEC dedicated SCFE Node cooperation Alarm to experts XEC PC 1 For expert use Important controls are implemented in SCS nodes Detector, storage, dewar, purifiers

  19. Temperature Sensor Liquid circulation Purifier Cartridge Molecular sieves, 13X 25g water Freq. Inverter OMRON • Circulate xenon in liquid phase • Circulation pump • 100liter/h@3175rpm, Dp = 0.2MPa • Molecular sieves • >24 g water absorption PT

  20. Succeeded! Circulation period 50.63Hz ! • Circulation speed evaluation • change of the surface level after closing the inlet valve • 3.6% / 30sec ~ 432% / h • 1% corresponds to 0.165 liter • 0.165x 432%/h = 71 liter /hour

  21. 205 h 2/Dec 180 h 23/Nov 70 h 14/Nov Liquid-phase purification • Light yield for 17.6MeV g saturated around 23/Nov (180h purification time) • Necessary to continue longer than we expected • 5 hours purification was enough in the LP test (100L LXe) • Probably due to • slower circulation speed (100L/h  70L/h) • Worse initial condition compared to the LP • Needed longer time to prepare monitoring tools due to PMT HV feedthru problem (reported later) • Noise from the pump (freq. inverter?) affected other detectors C-W run 17.6MeV gamma

  22. Absorption Length • Ratio Data/MC vs distance fitted with an exponential curve. • Inner and Outer face PMTs • Cosine of incidence angle < -0.2 • Slope compatible with zero(no absorption). l> 3 m @95 % C.L. After 50 h purification, 4/Nov Alpha source PMT

  23. 2D display, charge/time 2D hist, charge:PMT# 3D reconstructed position Performance Waveform 1D histograms Charge:event#

  24. PMT Calibration • PMT calibration • LED • PMT gain • Flushing LEDs at different intensities • Npe~1/s2 • Time offset calibration • Viewing one LED flushing with many PMTs simultaneously • Alpha • QE and absorption length evaluation • Liquid and cold gas

  25. Time offset determination • Possible method only in non-segmented detector like ours • c’ is obtained by using all data Different LEDs viewed by one PMT Measured Flashing Time Of i-th LED Speed of LED light Offset of j-th PMT Tj RD run (ultra low) ti –ti-1 1/c’ txe - tTC li,j-li-1,j

  26. C-W run • Li at 14.6, 17.6 MeV • B at 4.4, 11.7 and 16.1 MeV • Details in Giovanni’s presentation

  27. CEX process p-pp0n p0(28MeV/c)  g g 54.9 MeV < E(g) < 82.9 MeV LH2 target NaI tagging counter 3x3 crystals, APD readout Pb collimator panel in front of the Xe detector Eg 170o q Eg Eg p0 175o q 54.9MeV 82.9MeV 1.3MeV for q>170o 0.3MeV for q>175o Eg CEX run - Pi0 calibration g

  28. LH2 Target • Pressure test of cell 4.5 bar (abs) • Time to liquefy • 80 min from start of LHe flow • Liquid stability • 1.2 bar operating pressure (96% cell full) • 1.3% RMS, 6% max excursion • Liquid Helium consumption • 2.4%/h • 42h auto

  29. NaI Detector beam axis+/- 21 (+/- 0.07) deg. Up/down+/- 60 (+/- 0.005) deg.

  30. CEX run data analysis, preliminary • Position cut • Cut shallow events (< 2cm) • Select only center events ( |u|,|v| < 5cm) • Position correction • Pile-up rejection by light distribution • Select center event on NaI detector • Not applied QE correction • If applied worsen resolution. • Pedestal has 2% spread • Needs better baseline evaluation • Check hardware for 2008 run 55MeV gamma pedestal sup = 2.4% FWHM = 6.5%

  31. Practical resolution by tXe – tsci 280ps Time Measurement Intrinsic resolution by T-B analsysis • Using only 12 PMTs around the center 115ps • T : weighted mean of inner PMT timing after subtracting photon propagation time • Effective velocity 10cm/ns

  32. Position Measurement • Using collimator run data,

  33. Physics run • RD run gamma energy • Data : Xe self trigger threshold=3.5V • MC : RD event generation + event overlaps + trigger simulation • Vertical scale is scaled assuming, • Mu stop = 5e6 • Calorimeter acceptance = 0.1 • Calorimeter detection efficiency = 0.6

  34. Problems found in 2007andsolutions for 2008 run

  35. Feedthru • We could not apply required voltage on all PMTs at the beginning • We found that this had been caused by spark in the feedthru • Needed to prepare “new ones” for 2007 run • Commercial products or hand made?

  36. How did we make new ones? Wataru’s Design Air Metal body Glass insulator Xe • Production procedure • Fix pins in the holes and fill with silica • Bake in argon atmosphere • Cool down No need to change connector Replacement can be done quickly Body made of insulator (not metal)

  37. Installation • Oct. 10-12: LXe recovery to 1000L Dewar • Oct. 13-14: GXe recovery • Oct. 15-16: Mounting new flanges & testing • With flushing dry nitrogen gas in the cryostat • Oct 17-19: evacuation • Gas filling ~0.13MPa • Successfully took gas alpha data at 800, 900, 1000V • LED data • Liquid transfer started at almost same time • Until the detector is cooled we continued LED and alpha DAQ • 21, 22, 23/Oct transfer speed 15~20 liter/hour • Completed early morning of 24/Oct • 2 weeks interruption of DAQ

  38. PMT status after replacing feedthru • Stable operation after replacing feedthru • LED intensity optimization • Better gain evaluation than before

  39. As a bonus… • Xenon recovered through purifier • Practice to transfer the liquid to the dewar • Water contamination suppressed • Cryostat was warmed up to ~220K • 1st experience of temperature cycle • Test of the gasket • RGA: I = 3.8x10-13A  5.7x10-7 mbarl/sec 132Xe++

  40. Future plan (currently ongoing) • Replacement during winter shutdown • 48 pin x 6 x 4 • Need to change connectors • Used in the small prototype and PMT test chamber ceramic Kyocera Ultra High Vacuum Feedthrough welding

  41. PMTs ? Cables Pipes Heat load • Pressure is slowly increasing under normal operating condition. • Refrigerator cooling power: 200W • Expected redundant cooling power: ~100W • Calculation based on LP modeling • PMT:37W (Vmean = 775V) • Conduction:64W • Cable (50), Chimney (4.8), SI(3.1), Support(6.3) • Heat income through Cu cooling pipes was not taken into account • Cu heat conduction: 390 W/m/K • 10mm diam 1mmt pipe, 20cm • 390 x 135 (K) x 2.83x10-5/0.2 (m) = 7.4W • 6 pipes  44.4W • Cf. Steel heat conduction ~20W/m/K • There seems to be other heat leak… • Larger than 50W • Super-insulator?

  42. LN2 cooling pipe • Cu pipe feedthru  SUS pipe feedthru

  43. Cooling Power

  44. PT Blow up of low temperature valve • Low temperature valves blew up • A few liters of liquid xenon was lost • Purifier cryostat was opened • Misleading valve design • No documentation on the manual • Piping was modified and no valve is in use now Wilson Seal Air Wilson Seal xenon Plug or shaft Cup nut collar bellows O-ring Liquid xenon

  45. PMT Signal Splitter • Spark in feedthru’s destroyed protection zener diodes on PMT splitter boards • base-line shift at splitter output • Signal was out of range of WFD • Fixed by replacing all zener diodes DRS splitter PMT Trigger

  46. Light YieldDiscrepancy between a and g data • Energy scale discrepancy btw alpha and gamma • Too small light yield from g events (~1/2) • Not due to magnetic field • Confirmed by taking C-W data w/o COBRA field • Purity seems good • Improvement and plateau of light yield of both gammas and alphas • Have a look on WF a

  47. g a e Xe Xe e e Xe Xe Waveforms • triplet= 22 ns • recomb= 45 ns tg = 34 ns ta = 21 ns ! Careful treatment of electronics time constant is necessary Before purification Q/A was 1.93+/-0.02 in LP test A Q Electronegative impurity? Oxygen??

  48. Electronegative Impurity Removal • O2 getter cartridge • Developed for LAr use at CERN • be mounted at the exit MEG liquid-phase purifier with by-pass valves • Preparing an oxygen purity monitor also • will be ready at an early stage of 2008 run

  49. Current Status and Schedule • Xenon recovered to the 1000 liter dewar • Gas analysis will be done on site • Cryostat is opened now • All PMTs and cables are checked • Replacing feedthru is in progress • LN2 cooling pipe modification • Cryostat will be moved back to PiE5 at the end of March • Evacuation, liquid transfer, purification • Ready on 19/April • Schedule at http://meg.web.psi.ch/subprojects/install/xenon.html